OpenStudy (anonymous):

Determine whether the series converges, and if so find its sum. (i) \sum^{\infty}_{k=1}\frac{1}{9k^2+3k-2} (ii) \sum^{\infty}_{k=1}\frac{4^{k+2}}{7^{k-1}} (iii) \sum^{\infty}_{k=1}(\frac{1}{(k+p)(k+p+1)})

OpenStudy (anonymous):

can you post the real quest... i did not get it

OpenStudy (anonymous):

yup i understand the series look

OpenStudy (anonymous):

have u heard about telescopic series

OpenStudy (anonymous):

if so, you can apply it to find out it converges to a number

OpenStudy (perl):

(i) $$\sum^{\infty}_{k=1}\frac{1}{9k^2+3k-2}$$ (ii)$$ \sum^{\infty}_{k=1}\frac{4^{k+2}}{7^{k-1}}$$ (iii)$$ \sum^{\infty}_{k=1}(\frac{1}{(k+p)(k+p+1)})$$

OpenStudy (anonymous):

ops too many... am talking about the first

OpenStudy (anonymous):

thank you "Perl" . it seems that "Krishtika" is not listening :)

OpenStudy (michele_laino):

hint: we can rewrite the first sum as below: \[\Large \frac{1}{3}\sum\limits_{k = 1}^{ + \infty } {\left( {\frac{1}{{3k - 1}} - \frac{1}{{3k + 2}}} \right)} \]

OpenStudy (anonymous):

i'm back so u people got the solution

OpenStudy (thomas5267):

\[ \sum^{\infty}_{k=1}\frac{4^{k+2}}{7^{k-1}}=16\times7\sum^{\infty}_{k=1}\frac{4^{k}}{7^{k}} \] The third one looks convergent to me but I have no idea how to evaluate the sum.

OpenStudy (thomas5267):

Partial fraction decomposition on the third one just like the first one. This will yield a telescoping series.

OpenStudy (loser66):

for 3) make comparison test with \(\dfrac{1}{k^2}\), we can get the sum is convergent. to find the sum, use limit \[lim_{t\rightarrow \infty}\int_1^t \dfrac{1}{k+p}-\dfrac{1}{k+p+1}dk\] After calculating, I got 1 :)

OpenStudy (anonymous):

that method is an advanced one compared to Telescopic series partial fraction way....the second is much easier

OpenStudy (anonymous):

@Loser66, comparing the sum to the integral doesn't tell us much about the actual sum, but rather an upper (or lower, depending on the summand's behavior) bound.